Question 4. (20 pts) Use polar coordinates to find the volume of the solid between z=2? + y² and 2 = 3 – 22 – y?.
Answer #1

polar coordinates » 8² - x²rty where r is positive always Z 206+y? $ Z = 3-xx² - y² Z = 2 Z = 3-(x+y) Z: 3 r = ± 2 2 Ogre + 3